In Drosophila, mutations in the genes serrate (Ser), dachs (d+), and eyeless (ey+) leads to abnormal wings, short legs, and blind flies, respectively. A cross between wild type flies and flies that display the serrate,dachs, and eyeless phenotypes results in an F1 that have abnormal wings. What are the expected genotypes if we cross F1 flies? What are the expected phenotypes ratios?

Your peers decide to do the same exact F1 cross. Of a total of 1,000 flies, they obtain 430 serrate flies and, 138 serrate and eyeless flies. Based on the expected phenotype ratios you obtained above, perform a chi-square test to make sure your expected ratios are in accordance with the observed values.
Answer #1

Ser= abnormal wings, dachs= short legs, eyeless= blind flies, so the cross between Ss Dd Ee x ss dd ee = 1/2 x 1/2 x 1/2= 1/8 have abnormal wings in F1, so

S s
s Ss ss
s Ss ss

so, chances of mutated serated is 1/2 out of 4, similarly the other two

EXPECTED=serated = 1/2 OR 0.5 , ser and eyeless= 1/2x1/2= 1/4.= 0.25

OBSERVED= given ser =430/1000= 0.43; ser eyeless 138/1000= 0.138